OpenStudy (anonymous):

- Prove that 1+sin(x)- Cos(x)= 2 Sin(1/2x) [ Cos(1/2x) +Sin(1/2x) ] I'm stuck at the last step for like an hour or so reached this 1+sin(x)- Cos(x) = Sin(X) + 2Sin^2(1/2x) Help please :)

OpenStudy (kc_kennylau):

Welcome to OpenStudy :D You may want to read the code of conduct ( http://openstudy.com/code-of-conduct ). Don't forget to click on the button called "Best Answer" to give a medal to the person who helped you :)

OpenStudy (kc_kennylau):

Remember \(1−\cos(2x)\)?

OpenStudy (anonymous):

No ++"

OpenStudy (anonymous):

& yeah sure :) for the medal thing ^^

OpenStudy (kc_kennylau):

lol that's just my welcome msg for new ppl like u

OpenStudy (kc_kennylau):

ok if you look into that step, you'll find that \(\sin(x)\) exists in both sides

OpenStudy (anonymous):

yeah , the 1+sin(x)- Cos(x) = Sin(X) + 2Sin^2(1/2x) , right ?

OpenStudy (kc_kennylau):

so you just have to prove that \(1-\cos(x)=2\sin^2\left(\dfrac12x\right)\)

OpenStudy (anonymous):

I did it ! \[1- cosx = 2\sin ^{2 \left(\begin{matrix}1 \\ 2\end{matrix}\right)} x \] which is equal to \[1-\left( 1-2\sin ^{2\left(\begin{matrix}1 \\ 2\end{matrix}\right)} \right)x = 2\sin ^{2\left(\begin{matrix}1 \\ 1\end{matrix}\right)}x \] :D Thanks so much ! really appreciate it and for the welcome message too :D

OpenStudy (kc_kennylau):

you can use \frac12 for \(\frac12\) :)

OpenStudy (kc_kennylau):

and no problem :)

OpenStudy (anonymous):

yeah just noticed it :) Thanks for your help ^^